OpenStudy (anonymous):

y = -8/5x-5 how do I go about solving?

OpenStudy (dumbcow):

what are you solving for?

OpenStudy (anonymous):

it says to graph each line

OpenStudy (dumbcow):

every linear equation is of the form y = mx + b m = slope (rise/run) b = y-intercept (where it crosses y-axis) ->y = -8/5 x - 5 slope = -8/5 y-intercept = -5 Go to graph and put first point on y-intercept then use slope to find second point

OpenStudy (anonymous):

Thank you so much I thought I did it right the first time. LOL:) I appreciate your help!

OpenStudy (dumbcow):

your welcome